Students taking a module have a year mark above 75 with probability 20%, a year mark between

30 and 75 with probability 60%, and a year mark below 30 with probability 20%. A student with a

year mark above 75 passes the exam with probability 70%, a student with a year mark between 30

and 75 passes the exam with probability 65% and a student with a year mark below 30 passes the

exam with probability 15%.

(a) Construct a probability tree for the described scenario. (6)

(b) What is the probability that a randomly chosen student passes the exam? (6)

(c) If a student has a year mark below 75, what is the probability that the student passes the

exam? (4)



1
Expert's answer
2021-06-22T11:12:05-0400

(a) Let "A" denotes the event "a year mark is above 75".

Let "B" denotes the event "a year mark is between 30 and 75".

Let "C" denotes the event "a year mark is below 30 ".

Given "P(A)=0.2, P(B)=0.6, P(C)=0.2."


Let "Y" denotes the event "a student passes the exam".

Given "P(Y|A)=0.70, P(Y|B)=0.65, P(Y|C)=0.15."




(b) The Law of Total Probability


"P(Y)=P(A)P(Y|A)+P(B)P(Y|B)+P(C)P(Y|C)"

"=0.2(0.70)+0.6(0.65)+0.2(0.15)=0.56"

(c)


"P(Y|B\cup C)=P(B)P(Y|B)+P(C)P(Y|C)"

"=0.6(0.65)+0.2(0.15)=0.42"